Shayna owns a jewel store. She handed you a golden-colored cube for you to buy it. Out of curiosity for the element, you search on the internet about geology and lookup gold in the material table and read that its density is 19.3 g/cm3. You measure the cube and find that it is 2 cm on each side, and weighs 40 g. What is its density? Is it gold?


1
Expert's answer
2021-10-24T18:27:51-0400

1) Determine the volume:


"V=a^3=2^3=8\text{ cm}^3."

2) Determine density:


"\rho=m/V=5\text{ g/cm}^3."

This is not gold.